OpenStudy (blackstreet23):

in each part, compare the natural domains of f and g a. f(x) = ((x^2) + x) / (x+1) b. f (x) = ((x√ x) + √ x) / (x+1)

OpenStudy (blackstreet23):

Please explain me how to do it. I am totally lost

OpenStudy (triciaal):

a) look at the denominator (the divisor) x cannot be = -1 because you do not want to divide by zero look at the numerator x^2 + x can be factored as x(x + 1) if x = -1 then you would be multiplying by zero anything by zero = zero

OpenStudy (triciaal):

where is g in this problem?

OpenStudy (blackstreet23):

ohh yeah i firgot in each part, compare the natural domains of f and g a. f(x) = ((x^2) + x) / (x+1) ; g(x) = x b. f (x) = ((x√ x) + √ x) / (x+1) ; g(x) = √x